HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

MARVEL ANALYSIS OF EXPERIMENTAL TRANSITION DATA AND REFINEMENT OF THE^{32}S^{16}O SPECTROSCOPIC MODEL: THE IR/VIS

LINE LIST

\underline{RYAN P. BRADY}, SERGEI N. YURCHENKO, JONATHAN TENNYSON, Department of Physics and Astronomy, University College London, Gower Street,

WC1E 6BT London, United Kingdom.

We present a new rovibronic IR/Vis line list for the transient diatomic sulfur monoxide (^{32}S^{16}O) computed through our variational code Duo^{1} using a semi-empirical spectroscopic model consisting of potential energy curves, spin-orbit curves, electronic angular momentum curves, (transition) dipole moment curves as well as other couplings. The underlying ab initio spectroscopic model of SO was taken from Brady et al. (2022)^{2} which was refined by fitting to a comprehensive experimentally derived set of energies of SO. To this end, an experimental set of 50106 transitions, 49613 of those being non-redundant, have been compiled through the analysis of 29 experimental sources. A self-consistent set of 8850 rovibronic energy levels for the X ^{3}Σ^{−}, a ^{1}Δ, b ^{1}Σ^{+}, A ^{3}Π, B ^{3}Σ^{−}, and C ^{3}Π electronic states has been generated with the MARVEL algorithm^{3} covering rotational and vibrational quantum numbers J ≤ 69 and ν ≤ 30, respectively, and energies up to 52350.40 cm^{−1} (≥ 191 nm). A large gap in our network between 12300–20500 cm^{−1} exists due to lack of vibrational data for X ^{3}Σ^{−}, a ^{1}Δ, b ^{1}Σ^{+} with no coverage of the c ^{1}Σ^{−}, A ^{3′′}Σ^{+}, and A′ ^{3}Δ states. Our refined spectroscopic model represents the X ^{3}Σ^{−}, a ^{1}Δ, b ^{1}Σ^{+}, and A^{3}Π MARVEL energies to a weighted root-mean-square error of 4.0 × 10^{−5} cm^{−1}, 1.5 × 10^{−6} cm^{−1}, 2.7 × 10^{−7} cm^{−1}, and 4.8 × 10^{−3} cm^{−1}. We do not refine the UV energies here because of the many perturbations in the B ^{3}Σ^{−} and C ^{3}Π energies due to their overlapping potentials and large coupling. We do not have data connecting the B ^{3}Σ^{−}(ν = 0) state which is required to model the perturbations correctly, and so we leave this to a future project.

1doi:10.1016/j.cpc.2015.12.021S.N. Yurchenko, L. Lodi, J. Tennyson, and A.V. Stolyarov, Duo: A general program for calculating spectra of diatomic molecules, Comput. Phys. Commun., 202, 262 (2016), https://github.com/ExoMol/Duo

2doi:10.1039/D2CP03051AR. P. Brady, S. N. Yurchenko, G.-S. Kim, W. Somogyi and J. Tennyson, An ab initio study of the rovibronic spectrum of sulphur monoxide (SO): diabatic vs. adiabatic representation, Phys. Chem. Chem. Phys., 24, 24076 (2022)

3doi:10.1016/j.jms.2007.07.005T. Furtenbacher, A. G. Császár and J. Tennyson, MARVEL: measured active rotational-vibrational energy levels, J.Mol. Spectrosc., 245, 115-125 (2007)

p-number: p170  Submitted on Thu, 15 Jun 2023 11:46:48 +0200

p-number: p170  Submitted on Thu, 15 Jun 2023 11:46:48 +0200